The equation y= 5 sin (3x-4) where y is in millimetres,x is in metres and t is in seconds, represent a a wave motion. Determine
a) Frequency (b) Period (c) speed of the wave
1
Expert's answer
2020-10-21T09:52:18-0400

The wave equation can be written as follows:


"y=Asin(kx-\omega t),"

here, "A" is the amplitude of the wave, "k" is the wavenumber, "\omega" is the angular frequency of the wave.

By the definition of the question, we have:


"y=5sin(3x-4t)."

From this equation we can get the angular frequency and wavenumber:


"\omega=4\ s^{-1}, k=3\ m^{-1}."

a) We can find the frequency of the wave from the formula:


"f=\dfrac{\omega}{2\pi}=\dfrac{4\ s^{-1}}{2\pi} = 0.64\ Hz."

b) We can find the period of the wave from the formula:


"f=\dfrac{1}{T},"

here, "T" is the period of the wave.

Then, we get:


"T=\dfrac{1}{f}=\dfrac{1}{0.64\ Hz}=1.56\ s."

c) We can find the speed of the wave from the formula:


"v=\dfrac{\omega}{k}=\dfrac{4\ s^{-1}}{3\ m^{-1}}=1.33\ ms^{-1}."

Answer:

a) "f=0.64\ Hz."

b) "T=1.56\ s."

c) "v=1.33\ ms^{-1}."